# WavJEPA: Semantic learning unlocks robust audio foundation models for raw waveforms

### v1

@@ Abstract
Learning audio representations from raw waveforms overcomes key limitations of spectrogram-based audio representation learning, such as the long latency of spectrogram computation and the loss of phase information. Yet, while self-supervised speech representation learning from raw waveforms has been remarkably successful, these approaches have not achieved similar feats for general-purpose audio representation learning from waveforms. Here, we propose WavJEPA, a waveform-based version of the Joint-Embedding Predictive Architecture. WavJEPA leverages high-level semantic representation learning to tackle the shortcomings of representation learning at the speech unit or token level. We show that this approach substantially outperforms state-of-the-art time-domain audio foundation models across a wide variety of downstream benchmark tasks, while requiring considerably fewer computational resources. Additionally, to overcome the performance drop that time-domain models typically exhibit in noisy and reverberant real-world acoustic environments, we present WavJEPA-Nat. WavJEPA-Nat is a multi-channel extension of the WavJEPA architecture trained on simulated naturalistic scenes. We find that WavJEPA-Nat is highly robust to reverberation and noise. These results highlight the feasibility and computational efficiency of general-purpose audio representation learning from raw waveforms, showcasing the potential for low-latency, robust time-domain audio foundation models for real-world applications.
@@ 1 Introduction
State-of-the-art approaches for self-supervised general-purpose audio representation learning predominantly operate on spectrograms, that is, time-frequency representations of sound clips (Turian et al., 2022; Yadav et al., 2024; Chen et al., 2023; Gong et al., 2022; Yadav & Tan, 2024). However, these approaches suffer from two fundamental limitations: The latency introduced by the short-time Fourier transform (STFT) required for spectrogram computation impedes real-time deployment (Luo & Mesgarani, 2019), and (2) the loss of phase information reduces the performance on generative audio tasks (Luo & Mesgarani, 2019; Li et al., 2025). In contrast, time-domain models, which learn directly from raw audio waveforms, achieved remarkable success in speech representation learning (Baevski et al., 2020; Hsu et al., 2021; Chen et al., 2021). Crucially, end-to-end audio representation learning from raw waveforms overcomes the key limitations of spectrogram-based audio representation learning (long latency and loss of phase information) (Luo & Mesgarani, 2019). Yet, when state-of-the-art approaches for speech representation learning are trained for general-purpose audio representation learning, their performance is less strong (La Quatra et al., 2024). Furthermore, existing time-domain models exhibit significant degradation in noisy and reverberant acoustic environments compared to their spectrogram-based counterparts, limiting their effectiveness for real-world applications (Yuksel et al., 2025).
@@
To improve these shortcomings, we propose WavJEPA, a novel framework for end-to-end general-purpose audio representation learning from raw waveforms. The idea behind WavJEPA is that the semantic learning capabilities of joint-embedding predictive architectures (JEPAs) (Bardes et al., 2024; Assran et al., 2023; LeCun, 2022) can be leveraged to overcome the limitations of learning representations at the token or speech unit level, which is the typical approach of audio foundation models operating on raw waveforms. Instead, WavJEPA learns semantic representations by predicting the latent representations of training targets from a temporally distributed context representation of the same sound wave.
@@
WavJEPA is the first framework applying semantic learning to general-purpose audio representations in the time domain, surpassing state-of-the-art time-domain approaches on the HEAR (Turian et al., 2022) and ARCH (La Quatra et al., 2024) benchmark suites while requiring only a fraction of the computational resources. Additionally, we address the degraded performance of time-domain models in real-world sound scenes with WavJEPA-Nat, a multi-channel extension of the WavJEPA framework trained on simulated real-world sound scenes. Evaluation on Nat-HEAR (Yuksel et al., 2025), a naturalistic version of the HEAR benchmark suite, demonstrates that WavJEPA-Nat exceeds the robustness of other time-domain foundation models to noise and reverberation. We furthermore elucidate the critical factors for semantic representation learning from raw waveforms through extensive ablation studies, targeting context-target sampling, top-$K$ averaging and the optimal ratio between real-world scenes and dry sound clips. In sum, WavJEPA and WavJEPA-NAT demonstrate that robust time-domain approaches for audio representation learning are feasible and efficient, opening the door to low-latency audio foundation models for real-world applications.
@@ 2 Related work
Spectrogram-based audio representation learning: These approaches aim to learn general-purpose representations from the time-frequency representation of a sound clip (spectrogram) calculated with a short-time Fourier transform. Masked auto-encoder (MAE) approaches achieve state-of-the-art performance on benchmark suites, learning rich audio representations by reconstructing masked spectrogram patches (He et al., 2022; Yadav et al., 2024; Gong et al., 2022; Chong et al., 2023; Baade et al., 2022; Huang et al., 2022; Yadav & Tan, 2024). Other approaches – inspired by the success in the visual domain (Grill et al., 2020; Bardes et al., 2024; Assran et al., 2023) – avoid reconstructing the original spectrogram input space, instead predicting targets in the latent space (Niizumi et al., 2023; Fei et al., 2024; Chen et al., 2023; 2024).
Waveform-based audio representation learning: Representation learning from raw waveforms is based on predictive or contrastive self-supervised learning strategies at the token or speech unit level (Baevski et al., 2020; Hsu et al., 2021; Chen et al., 2021). More recently, Data2Vec (Baevski et al., 2022) introduced a modality-agnostic framework for training across speech, vision, and text domains, leveraging a teacher-student approach. They demonstrated that the proposed latent prediction framework achieves state-of-the art on speech recognition with minimal fine-tuning. While these frameworks have proven extremely fruitful for speech representation learning, they have been less successful in learning general-purpose audio representations (Yadav et al., 2024; Turian et al., 2022; Yuksel et al., 2025; La Quatra et al., 2024).
Representation learning with joint embedding predictive architectures (JEPAs): Recent work demonstrated that JEPA models efficiently learn semantic image representations by predicting latent representations of parts of the input image (that is, training targets) from a context representation of other parts of that same image (Assran et al., 2023; Bardes et al., 2024). Based on this success, others applied JEPA models to spectrograms (Fei et al., 2024), EEG signals (Guetschel et al., 2024) and fMRI measurements (Dong et al., 2024), highlighting the versatility of the JEPA framework.
@@ 3.1 The WavJEPA framework
Our proposed architecture and approach for learning general-purpose audio representations from raw waveforms are illustrated in Figure 1. The WavJEPA framework comprises a waveform encoder, context encoder, target encoder and a predictor. WavJEPA’s objective is to predict latent representation of various targets blocks based on a single context block extracted from the same sound wave. As waveform encoder, we use the feature encoder of Wav2Vec 2.0, which is composed of stacked temporal convolution layers (Baevski et al., 2020). Similar to the original I-JEPA architecture (Assran et al., 2023), a Vision Transformer (ViT) (Dosovitskiy et al., 2021) is used for the target encoder, context encoder and predictor. Detailed specifications of the framework components can be found in Appendix A. In the following, we describe the main components of the WavJEPA framework.
@@
Target encoder and learning objective: In this waveform-based approach, latent representations of the sound wave embeddings constitute the targets. The target encoder $E(\cdot)$ converts the whole waveform embedding $w$ into a latent target representation. Similar to Baevski et al. (2022; 2020), the outputs of the top $K$ layers are instance-normalized (Ulyanov et al., 2017) and averaged. For each time step $i\in[1\dots N]$, we obtain a target embedding $y_{i}\in\mathbb{R}^{768}$. For each target block $k\in[1\dots K_{target}]$, we select the tokens $y^{k}=\{y_{t^{k}_{1}},...,y_{t^{k}_{M_{target}}}\}$ corresponding to the target block indices $t_{1}^{k},\dots,t^{k}_{M_{target}}$ and compute the $L_{2}$ distance between the predicted target representation $\hat{y}^{k}$ and the actual training target $y^{k}$. The final loss corresponds to the average error across targets.
@@ 3.2 Experimental set-up WavJEPA
We train WavJEPA on the unbalanced training set of AudioSet, which consists of 1.74 million 10-second sound clips scraped from YouTube (Gemmeke et al., 2017). Each sound clip was resampled to 16 kHz and mean centered to enforce equal loudness across sound clips. We then randomly sampled 8 sections of 2 s from each sound clip, effectively increasing the batch size by a factor of 8 in a computationally efficient manner. Finally, each instance is instance normalized (Ulyanov et al., 2017). The waveform encoder converts each 2 s instance into an embedding $w^{200\times 768}$, effectively resampling the audio to 100 Hz with a stride of 10 ms and a receptive field size of 12.5 ms.
@@
We trained WavJEPA for 375,000 steps using a batch size of 32 on two NVIDIA H100 94 GB GPUs. Given our in-batch sampling factor of 8, we boost our effective batch size to 256. We use the AdamW optimizer (Loshchilov & Hutter, 2019) with a weight decay coefficient $\lambda_{w}$ = 0.04. The learning rate schedule follows a cosine decay with linear warm-up over 100,000 steps, reaching a peak learning rate of $2\times 10^{-4}$ before decaying to zero.
@@ 3.3 The WavJEPA-Nat Framework
Data and sound wave embeddings: We use the pipeline of Yuksel et al. (2025) to transform AudioSet sound clips into naturalistic, spatialized sound scenes with reverberation and noise. In brief, we simulate naturalistic, spatialized scenes by using the room impulse response (RIR) simulator and binaural renderer provided by Soundspaces 2.0 (Chen et al., 2022), resulting in two-channel sound clips containing naturalistic spatial cues. To each sound scene, we add similarly spatialized noise clips from the WHAMR! database (Maciejewski et al., 2020). A full description of the sound scene generation can be found in Appendix D.
@@ 3.4 Downstream evaluation
Downstream tasks: We evaluated WavJEPA and WavJEPA-Nat on two large benchmark task suites for the evaluation of general-purpose audio foundation models: HEAR (Turian et al., 2022) and ARCH (La Quatra et al., 2024). We use the same subset of HEAR benchmark tasks as previously used in Yadav et al. (2024) but added DCASE2016 Task 2 (Mesaros et al., 2018b) as a time stamp-based task to evaluate the audio scene analysis capabilities of the models more in-depth. HEAR and ARCH contain a selection of complementary tasks and datasets for acoustic events and scene analysis, speech, and music. For more detailed description of tasks please see Appendix C.
@@
We additionally evaluated models on NatHEAR (Yuksel et al., 2025), a naturalistic version of the HEAR benchmark suite comprising high-quality simulations of real-world sound scenes with reverberation and noise, spatialized in two formats (either binaural and ambisonics). To accommodate the input format of single-channel models, we utilized the first channel (that is, the omndirectional microphone) of NatHEAR in the Ambisonics format (Zotter & Frank, 2019). For the dual waveform encoder approach of WavJEPANat, we used both channels of NatHEAR in a binaural format.
@@
Model fine-tuning for downstream evaluation: For the downstream evaluation on HEAR and ARCH benchmark tasks, we trained a shallow downstream classifier on representations that were extracted after self-supervised pre-training, following the exact fine-tuning procedures detailed by HEAR (Turian et al., 2022) and ARCH (La Quatra et al., 2024). Model weights were frozen after pre-training. Note that the difference between the fine-tuning approaches in HEAR and ARCH causes the differences in performance for tasks that are in both suites, for example, ESC50. Further, to evaluate WavJEPANat on HEAR, we duplicated the single-channel audio recordings of the original HEAR to make the input compatible with the dual waveform encoder architecture of WavJEPANat.
@@
Down stream evaluation metric $s(m)$: As the tasks in HEAR and ARCH vary considerably in terms of evaluation criteria and difficulty level, we calculate for each model $m$ a generalizability metric $s(m)$ to give an impression of the overall performance of a model, similar to Yang et al. (2021).This metric effectively ranks models as a function of the maximum improvement they obtain over the baseline model, normalized by the difference in scores between SOTA and the baseline for the specific task (see Appendix B). The baseline used here is HEAR-Naive, consisting of mel-spectrogram representations. For calculating this score, we included all the evaluated methods in all upcoming sections, including ablations.
@@
Model comparison: We compare the performance of WavJEPA to state-of-the-art self-supervised models using transformer architectures for representation learning from raw waveforms. We include Wav2Vec2.0 (Baevski et al., 2020), HuBERT (Hsu et al., 2021), WavLM (Chen et al., 2021), Data2Vec (Baevski et al., 2022), all pre-trained on large quantities of speech data. We furthermore include the recently released versions of Wav2Vec2.0 and HuBERT pre-trained on AudioSet (La Quatra et al., 2024) to assess their ability to learn general-purpose audio representations. For all models, we include both the Base (approximately 90 m parameters) and the Large version (300 m parameters). In comparison, WavJEPA has 90 m parameters.
@@ 3.5 Ablations:
To identify the critical parameters for a successful learning of general-purpose audio representations with the WavJEPA framework, we conducted comprehensive ablation studies on the pre-training parameters. Specifically, we examined the effect of sampling parameters for target and context blocks ($p_{target}$, $M_{context}$ and $M_{target}$) and the effectiveness of top-$K$ layer averaging for training targets. For WavJEPANat, we systematically assessed the impact of the ratio between clean and naturalistic sound scenes in the pre-training data. For all ablation studies, pre-training and downstream evaluation settings were similar to those of WavJEPA and WavJEPANat.

### v3

@@ Abstract
Learning audio representations from raw waveforms overcomes key limitations of spectrogram-based audio representation learning, such as the long latency of spectrogram computation and the loss of phase information. Yet, while self-supervised speech representation learning from raw waveforms has been remarkably successful, these approaches have not achieved similar feats for general-purpose audio representation learning from waveforms. Here, we propose WavJEPA, a waveform-based version of the Joint-Embedding Predictive Architecture. WavJEPA leverages high-level semantic representation learning to tackle the shortcomings of representation learning at the speech unit or token level. We show that this approach substantially outperforms state-of-the-art time-domain audio foundation models across a wide variety of downstream benchmark tasks, while requiring considerably fewer computational resources. Additionally, to overcome the performance drop that time-domain models typically exhibit in noisy and reverberant real-world acoustic environments, we present WavJEPA-Nat. WavJEPA-Nat is a multi-channel extension of the WavJEPA architecture trained on simulated naturalistic scenes. We find that WavJEPA-Nat is highly robust to reverberation and noise. These results highlight the feasibility and computational efficiency of general-purpose audio representation learning from raw waveforms, showcasing the potential for low-latency, robust time-domain audio foundation models for real-world applications.  [^1] All code and materials are available on https://github.com/labhamlet, and https://huggingface.co/labhamlet
@@ 1 Introduction
State-of-the-art approaches for self-supervised general-purpose audio representation learning predominantly operate on spectrograms, that is, time-frequency representations of sound clips (hear; mwmae; beats; gong2022ssast; ssam). However, these approaches suffer from two fundamental limitations: The latency introduced by the short-time Fourier transform (STFT) required for spectrogram computation impedes real-time deployment (conv-tas-net), and (2) the loss of phase information reduces the performance on generative audio tasks (conv-tas-net; li2025advances). In contrast, time-domain models, which learn directly from raw audio waveforms, achieved remarkable success in speech representation learning (wav2vec2; hubert; wavlm). Crucially, end-to-end audio representation learning from raw waveforms overcomes the key limitations of spectrogram-based audio representation learning (long latency and loss of phase information) (conv-tas-net). Yet, when state-of-the-art approaches for speech representation learning are trained for general-purpose audio representation learning, their performance is less strong (arch). Furthermore, existing time-domain models exhibit significant degradation in noisy and reverberant acoustic environments compared to their spectrogram-based counterparts, limiting their effectiveness for real-world applications (gram).
@@
To improve these shortcomings, we propose WavJEPA, a novel framework for end-to-end general-purpose audio representation learning from raw waveforms. The idea behind WavJEPA is that the semantic learning capabilities of joint-embedding predictive architectures (JEPAs) (vjepa; ijepa; LeCun2022APT) can be leveraged to overcome the limitations of learning representations at the token or speech unit level, which is the typical approach of audio foundation models operating on raw waveforms. Instead, WavJEPA learns semantic representations by predicting the latent representations of training targets from a temporally distributed context representation of the same sound wave.
@@
WavJEPA is the first framework applying semantic learning to general-purpose audio representations in the time domain, surpassing state-of-the-art time-domain approaches on the HEAR (hear) and ARCH (arch) benchmark suites while requiring only a fraction of the computational resources. Additionally, we address the degraded performance of time-domain models in real-world sound scenes with WavJEPA-Nat, a multi-channel extension of the WavJEPA framework trained on simulated real-world sound scenes. Evaluation on Nat-HEAR (gram), a naturalistic version of the HEAR benchmark suite, demonstrates that WavJEPA-Nat exceeds the robustness of other time-domain foundation models to noise and reverberation. We furthermore elucidate the critical factors for semantic representation learning from raw waveforms through extensive ablation studies, targeting context-target sampling, top-$K$ averaging and the optimal ratio between real-world scenes and dry sound clips. In sum, WavJEPA and WavJEPA-NAT demonstrate that robust time-domain approaches for audio representation learning are feasible and efficient, opening the door to low-latency audio foundation models for real-world applications.
@@ 2 Related work
Spectrogram-based audio representation learning: These approaches aim to learn general-purpose representations from the time-frequency representation of a sound clip (spectrogram) calculated with a short-time Fourier transform. Masked auto-encoder (MAE) approaches achieve state-of-the-art performance on benchmark suites, learning rich audio representations by reconstructing masked spectrogram patches (mae_approach; mwmae; gong2022ssast; MaskSpec; mae-ast; audiomae; ssam). Other approaches – inspired by the success in the visual domain (byol; vjepa; ijepa) – avoid reconstructing the original spectrogram input space, instead predicting targets in the latent space (byol-a; ajepa; beats; eat).
Waveform-based audio representation learning: Representation learning from raw waveforms is based on predictive or contrastive self-supervised learning strategies at the token or speech unit level (wav2vec2; hubert; wavlm). More recently, Data2Vec (data2vec) introduced a modality-agnostic framework for training across speech, vision, and text domains, leveraging a teacher-student approach. They demonstrated that the proposed latent prediction framework achieves state-of-the art on speech recognition with minimal fine-tuning. While these frameworks have proven extremely fruitful for speech representation learning, they have been less successful in learning general-purpose audio representations (mwmae; hear; gram; arch).
Representation learning with joint embedding predictive architectures (JEPAs): Recent work demonstrated that JEPA models efficiently learn semantic image representations by predicting latent representations of parts of the input image (that is, training targets) from a context representation of other parts of that same image (ijepa; vjepa). Based on this success, others applied JEPA models to spectrograms (ajepa), EEG signals (sjepa) and fMRI measurements (brainjepa), highlighting the versatility of the JEPA framework.
@@ 3.1 The WavJEPA framework
Our proposed architecture and approach for learning general-purpose audio representations from raw waveforms are illustrated in Figure 1. The WavJEPA framework comprises a waveform encoder, context encoder, target encoder and a predictor. WavJEPA’s objective is to predict latent representation of various targets blocks based on a single context block extracted from the same sound wave. As waveform encoder, we use the feature encoder of Wav2Vec 2.0, which is composed of stacked temporal convolution layers (wav2vec2). Similar to the original I-JEPA architecture (ijepa), a Vision Transformer (ViT) (dosovitskiy2020vit) is used for the target encoder, context encoder and predictor. Detailed specifications of the framework components can be found in Appendix A. In the following, we describe the main components of the WavJEPA framework.
@@
Target encoder and learning objective: In this waveform-based approach, latent representations of the sound wave embeddings constitute the targets. The target encoder $E(\cdot)$ converts the whole waveform embedding $w$ into a latent target representation. Similar to data2vec; wav2vec2, the outputs of the top $K$ layers are instance-normalized (instancenorm) and averaged. For each time step $i\in[1\dots N]$, we obtain a target embedding $y_{i}\in\mathbb{R}^{768}$. For each target block $k\in[1\dots K_{target}]$, we select the tokens $y^{k}=\{y_{t^{k}_{1}},...,y_{t^{k}_{M_{target}}}\}$ corresponding to the target block indices $t_{1}^{k},\dots,t^{k}_{M_{target}}$ and compute the $L_{2}$ distance between the predicted target representation $\hat{y}^{k}$ and the actual training target $y^{k}$. The final loss corresponds to the average error across targets.
@@ 3.2 Experimental set-up WavJEPA
We train WavJEPA on the unbalanced training set of AudioSet, which consists of 1.74 million 10-second sound clips scraped from YouTube (audioset). Each sound clip was resampled to 16 kHz and mean centered to enforce equal loudness across sound clips. We then randomly sampled 8 sections of 2 s from each sound clip, effectively increasing the batch size by a factor of 8 in a computationally efficient manner. Finally, each instance is instance normalized (instancenorm). The waveform encoder converts each 2 s instance into an embedding $w^{200\times 768}$, effectively resampling the audio to 100 Hz with a stride of 10 ms and a receptive field size of 12.5 ms.
@@
We trained WavJEPA for 375,000 steps using a batch size of 32 on two NVIDIA H100 94 GB GPUs. Given our in-batch sampling factor of 8, we boost our effective batch size to 256. We use the AdamW optimizer (adamw) with a weight decay coefficient $\lambda_{w}$ = 0.04. The learning rate schedule follows a cosine decay with linear warm-up over 100,000 steps, reaching a peak learning rate of $2\times 10^{-4}$ before decaying to zero.
@@ 3.3 The WavJEPA-Nat Framework
Data and sound wave embeddings: We use the pipeline of gram to transform AudioSet sound clips into naturalistic, spatialized sound scenes with reverberation and noise. In brief, we simulate naturalistic, spatialized scenes by using the room impulse response (RIR) simulator and binaural renderer provided by Soundspaces 2.0 (chen22soundspaces2), resulting in two-channel sound clips containing naturalistic spatial cues. To each sound scene, we add similarly spatialized noise clips from the WHAMR! database (Maciejewski2020WHAMR). A full description of the sound scene generation can be found in Appendix D.
@@ 3.4 Downstream evaluation
Downstream tasks: We evaluated WavJEPA and WavJEPA-Nat on two large benchmark task suites for the evaluation of general-purpose audio foundation models: HEAR (hear) and ARCH (arch). We use the same subset of HEAR benchmark tasks as previously used in mwmae but added DCASE2016 Task 2 (dcase) as a time stamp-based task to evaluate the audio scene analysis capabilities of the models more in-depth. HEAR and ARCH contain a selection of complementary tasks and datasets for acoustic events and scene analysis, speech, and music. For more detailed description of tasks please see Appendix C.
@@
We additionally evaluated models on NatHEAR (gram), a naturalistic version of the HEAR benchmark suite comprising high-quality simulations of real-world sound scenes with reverberation and noise, spatialized in two formats (either binaural and ambisonics). To accommodate the input format of single-channel models, we utilized the first channel (that is, the omndirectional microphone) of NatHEAR in the Ambisonics format (ambisonics). For the dual waveform encoder approach of WavJEPA-Nat, we used both channels of NatHEAR in a binaural format.
@@
Model fine-tuning for downstream evaluation: For the downstream evaluation on HEAR and ARCH benchmark tasks, we trained a shallow downstream classifier on representations that were extracted after self-supervised pre-training, following the exact fine-tuning procedures detailed by HEAR (hear) and ARCH (arch). Model weights were frozen after pre-training. Note that the difference between the fine-tuning approaches in HEAR and ARCH causes the differences in performance for tasks that are in both suites, for example, ESC50. Further, to evaluate WavJEPA-Nat on HEAR, we duplicated the single-channel audio recordings of the original HEAR to make the input compatible with the dual waveform encoder architecture of WavJEPA-Nat.
@@
Down stream evaluation metric $s(m)$: As the tasks in HEAR and ARCH vary considerably in terms of evaluation criteria and difficulty level, we calculate for each model $m$ a generalizability metric $s(m)$ to give an impression of the overall performance of a model, similar to superb.This metric effectively ranks models as a function of the maximum improvement they obtain over the baseline model, normalized by the difference in scores between SOTA and the baseline for the specific task (see Appendix B). The baseline used here is HEAR-Naive, consisting of mel-spectrogram representations. For calculating this score, we included all the evaluated methods in all upcoming sections, including ablations.
@@
Model comparison: We compare the performance of WavJEPA to state-of-the-art self-supervised models using transformer architectures for representation learning from raw waveforms. We include Wav2Vec2.0 (wav2vec2), HuBERT (hubert), WavLM (wavlm), Data2Vec (data2vec), all pre-trained on large quantities of speech data. We furthermore include the recently released versions of Wav2Vec2.0 and HuBERT pre-trained on AudioSet (arch) to assess their ability to learn general-purpose audio representations. For all models, we include both the Base (approximately 90 m parameters) and the Large version (300 m parameters). In comparison, WavJEPA has 90 m parameters.
@@ 3.5 Ablations:
To identify the critical parameters for a successful learning of general-purpose audio representations with the WavJEPA framework, we conducted comprehensive ablation studies on the pre-training parameters. Specifically, we examined the effect of sampling parameters for target and context blocks ($p_{target}$, $M_{context}$ and $M_{target}$) and the effectiveness of top-$K$ layer averaging for training targets. For WavJEPA-Nat, we systematically assessed the impact of the ratio between clean and naturalistic sound scenes in the pre-training data. For all ablation studies, pre-training and downstream evaluation settings were similar to those of WavJEPA and WavJEPA-Nat.